# SSUMamba: Spatial-Spectral Selective State Space Model for Hyperspectral Image Denoising

### v3

@@ III-D Architecture Design Details
IV Experiments
In this section, we demonstrate the denoising performance of the proposed SSUMamba through synthetic and real-world experiments.
IV-A Experimental Settings
IV-A1 Training and Testing Setting
@@ V Conclusion
This paper introduces a SSUMamba for HSI denoising. We incorporate the SSM with linear complexity to enable the global spatial-spectral correlation modeling in HSIs. By utilizing the Spatial-Spectral Alternating Scan Mamba block, our SSUMamba can effectively exploit the information flow in all directions in the HSI data and enhance the 3-D characteristic. Experimental results demonstrate that our proposed method achieves state-of-the-art performance.

### v4

@@ III-D Architecture Design Details
IV Experiments
In this section, we demonstrate the denoising performance of the proposed SSUMamba through synthetic and real-world experiments. Additionally, we conduct further ablation experiments to analyze the performance of each module and varying widths within the network.
IV-A Experimental Settings
IV-A1 Training and Testing Setting
@@ IV-C2 Earth Observing-1 HSI
IV-D Ablation Study
In this section, we present the performance of SSUMamba with varying widths to illustrate its performance under different number of parameters. Additionally, we conduct ablation experiments on different modules to demonstrate their individual contributions to performance.
| SSAS | Bidirectional | PSNR | SSIM | SAM |
| ✓ | ✓ | 34.55 | .9444 | .1047 |
| ✗ | ✓ | 34.40 | .9444 | .1056 |
| ✓ | ✗ | 34.26 | .9417 | .1086 |
IV-D1 Effectiveness of Different Modules
We assess the efficacy of the components introduced in SSUMamba, which include the SSAS and bidirectional Mamba. The model was trained on the ICVL training set and tested on the Houston 2018 HSI dataset with mixed noise. The results are presented in Table IV. The model lacking SSAS will consistently scan the HSI with a fixed priority, . Incorporating SSAS does not introduce additional parameters or computational complexity; rather, it aids the model in exploiting information flows the 3-D nature of the HSI data in all directions. This enhancement is reflected in improved PSNR, SAM, and nearly identical SSIM results. Furthermore, the introduction of the bidirectional Mamba further refines the global spatial-spectral correlation modeling, resulting in enhanced performance across all metrics.
IV-D2 Effect of Width
We further investigate the effectiveness of network width in our SSUMamba framework. To assess the impact of network width on performance, we conducted experiments with varying numbers of channels, which play a pivotal role in the network width. In the 6 SSAS Mamba blocks, the channel configuration is set as [], where means base channels. We compare the performance of SSUMamba with different numbers of base channels, including 32, 24, 16, and 12. We also compare the results with transformer-based methods, TRQ3D and SST. The model was trained on the ICVL training set and tested on the Houston 2018 HSI dataset with mixed noise. The results are presented in Table V. As we can see, the performance of our proposed SSUMamba continues to extend as the width is enhanced. Compared with transformer-based methods, which suffer from the complexity, our SSUMamba achieves better performance with fewer parameters and lower memory consumption. This is achieved thanks to the linear complexity facilitated by SSM and global spatial-spectral correlation modeling in our SSAS Mamba block.
| Channels | #Param. | Memory | PSNR | SSIM | SAM |
| 32 | 10.4M | 1,802M | 34.55 | .9444 | .1047 |
| 24 | 5.9M | 1,386M | 34.41 | .9413 | .1054 |
| 16 | 2.6M | 991M | 33.72 | .9373 | .1128 |
| 12 | 1.5M | 785M | 33.39 | .9287 | .1174 |
| Methods | #Param. | Memory | PSNR | SSIM | SAM |
| TRQ3D | 0.68M | 675M | 32.55 | .9194 | .1241 |
| SST | 4.14M | 1,600M | 31.07 | .9166 | .1390 |
@@ V Conclusion
This paper introduces a SSUMamba for HSI denoising. We incorporate the SSM with linear complexity to enable global spatial-spectral correlation modeling in HSIs. By utilizing the Spatial-Spectral Alternating Scan Mamba block, our SSUMamba can effectively exploit the information flow in all directions in the HSI data and enhance the 3-D characteristic. Experimental results with compared methods demonstrate that our proposed method achieves state-of-the-art performance. The ablation study further validates the effectiveness of the introduced components.